# Data Release 1 of the Dark Energy Spectroscopic Instrument

### v1

@@ Abstract
1 Introduction
1.1 The First Year of DESI Data
Elucidating the nature of dark energy and the physical mechanisms responsible for the accelerating expansion of the universe is one of the most important unsolved problems in physics and, arguably, all of science (Albrecht et al., 2006; Weinberg et al., 2013). To tackle this question, in 2021 May the Dark Energy Spectroscopic Instrument (DESI) collaboration began a 5-year survey to produce the most detailed three-dimensional map of the universe ever assembled (Levi et al., 2013; DESI Collaboration et al., 2016a, b). By measuring the baryon acoustic oscillation (BAO) feature and redshift space distortions at multiple cosmological epochs, DESI aims to place unprecedented constraints on the equation of state of dark energy, the gravitationally driven growth of large-scale structure, and the sum of the neutrino masses, as well as to explore the observational signatures of primordial inflation (DESI Collaboration et al., 2024a). Concurrently, DESI is carrying out an ambitious survey of stars in the halo of the Milky Way Galaxy in order to constrain the geometry, properties, and accretion history of its stellar halo, disk, and dark-matter halo (Cooper et al., 2023). Notably, DESI has the distinction of being the first Stage-IV dark-energy experiment to begin science operations (Albrecht et al., 2006).
@@ 1.1 The First Year of DESI Data
DESI identifies its primary spectroscopic targets using 14,000 deg2 of broadband optical and mid-infrared photometry from Data Release 9 (DR9) of the DESI Legacy Imaging Surveys111https://www.legacysurvey.org/dr9 (Zou et al. 2017; Dey et al. 2019; Moustakas et al. 2023; Myers et al. 2023a; Schlegel et al. 2025, in preparation; hereafter, the Legacy Surveys), and, in some cases, using Gaia stellar photometry (Gaia Collaboration et al., 2016). Specifically, DESI targets five broad classes of objects scaffolded in redshift: Milky Way survey (MWS) and backup program stars (Cooper et al. 2023; Dey et al. 2025, in preparation); bright galaxy survey (BGS) galaxies (; Ruiz-Macias et al., 2020; Hahn et al., 2023); luminous red galaxies (LRGs, ; Zhou et al., 2020, 2023); emission-line galaxies (ELGs, ; Raichoor et al., 2020, 2023); and QSOs (; Yèche et al., 2020; Chaussidon et al., 2023). For its cosmological analyses, DESI further differentiates QSO targets into “tracer” QSOs at lower redshift, , from Ly forest QSOs at , because above DESI uses the Ly forest as an independent probe of the matter-density field (DESI Collaboration et al., 2016a).
@@
After first light and a brief commissioning period, in 2020 December DESI began a five-month survey validation (SV) program which was designed to test the performance of the instrument and all its subsystems, and to validate its target-selection algorithms, data-reduction pipeline, and scientific deliverables (DESI Collaboration et al., 2024a). SV consisted of three successive phases: Target Selection Validation (SV1); Operations Development (SV2); and the One-Percent Survey (SV3), the last of which covered roughly 1% of the final 14,000 deg2 DESI footprint but to higher target completeness (DESI Collaboration et al., 2024a). These observations, particularly the One-Percent Survey data, resulted in a flurry of scientific activity222https://data.desi.lbl.gov/doc/papers/edr as well as DESI’s first measurement of the BAO peak in the BGS, LRG, ELG, and QSO 2-point correlation functions (Moon et al., 2023) and in the auto-correlation function of the Ly forest (Gordon et al., 2023). Subsequently, on 2021 May 14, DESI launched its 5-year main survey.333In detail, a small number of SV observations continued until 2021 June 10, but those data were acquired independently of the main survey and are included in the SV dataset (DESI Collaboration et al., 2024b).
@@
All the data obtained during SV were first made public on 2023 June 13 as part of the DESI Early Data Release (EDR; DESI Collaboration et al., 2024a, b).444https://data.desi.lbl.gov/doc/releases/edr Here, we present DESI Data Release 1 (DR1), which includes all the data obtained by DESI during its first 13 months of science operations (2021 May 14 through 2022 June 13) as well as a uniform reprocessing of all the SV data. Overall, DESI has been efficient, running well-ahead of schedule and occasionally obtaining redshifts for more than one million unique targets in a single month (Schlafly et al., 2023). In the first year of the main survey, DESI has measured confident redshifts for approximately 18.7M unique targets over more than deg2, including 13.1M galaxies, 1.6M quasars, and 4M stars, making DESI the largest extragalactic spectroscopic redshift survey ever conducted. For comparison, all five generations of the Sloan Digital Sky Survey (SDSS I/II, III, IV, and V; York et al., 2000; Dawson et al., 2013; Majewski et al., 2017; Blanton et al., 2017; Kollmeier et al., 2017), spanning 18 public data releases and approximately 25 years of operations,555https://www.sdss.org have cumulatively resulted in approximately 4M unique extragalactic spectra, making DESI DR1 nearly a factor of four larger than all previous SDSS programs combined.
@@
In §1.2 we outline the organization and contents of this paper and describe several important terms and concepts which will help orient readers interested in using DESI DR1. Complementarily, we recommend that readers interested in the cosmological results from DESI DR1 begin at the DESI portal666https://data.desi.lbl.gov/doc/papers/#year-1-cosmology-results for a high-level overview of the results before consulting the individual key papers (DESI Collaboration et al., 2024c, d, 2025a, e, 2025b, f).
1.2 A High-Level Guide to DESI and Data Release 1
@@
All DESI observations are made in the context of a survey and a program (Schlafly et al., 2023). As introduced in §1.1, DESI’s primary, 5-year scientific program is its main survey, which was preceded by three independent, successive surveys (SV1, SV2, and SV3), collectively called SV. In addition, an object can be observed as part of the special survey, which includes a miscellaneous assortment of targets and observations which are kept separate from the main survey (see §2.3.3).
Within each survey, objects are observed in one of three possible programs according to the acceptable lunar phase and on-sky observing conditions (see §2.3.1): bright, dark, or backup.777A fourth program name, other, is occassionally used for ad hoc or bespoke observations which do not easily fit into the three standard program names. By design, LRG, ELG, and QSO targets are assigned to the dark program and BGS and MWS targets are assigned to the bright program; meanwhile, the backup program consists entirely of stellar (Milky Way) targets, and special (secondary and tertiary targets) can be assigned to any program (see §2.2; Schlafly et al., 2023; Myers et al., 2023a; DESI Collaboration et al., 2024a). In general, objects within each survey/program combination are always treated independently, from target selection (§2.2) all the way through to spectral coaddition and redshift estimation (§3.1); in other words, the same astrophysical object can appear in two (or more) survey/program combinations. This strict separation of targets within each survey and program is a critical part of the large-scale structure pipeline which enables DESI to make precise cosmological measurements (Ross et al., 2025).
@@
| ProductionaaWithin the DESI data model, a spectroscopic production is referred to as a specprod. Each specprod corresponds to a uniform processing of the data using a well-defined and reproducible set of input files and software tags. | Stack | SurveysbbSee §2.3 for a description of the SV and main surveys. | RedshiftsccApproximate total number of unique redshifts summed over all surveys and programs. | Date |
@@
The primary spectroscopic production for DR1 is Iron, named after the Iron Mountain in Utah, USA; the Iron spectroscopic production includes data from the first 13 months of the DESI main survey, as well as a uniform reprocessing of the SV data which was previously released in the EDR (DESI Collaboration et al., 2024b). DR1 also includes a supplemental or ancillary spectroscopic production called Guadalupe (after the Guadalupe Mountains in Texas, USA), which includes just the first two months of the DESI main survey data (see Appendix A). We include Guadalupe in DR1 because it was used in some early DESI analyses (e.g., Moon et al., 2023; Gordon et al., 2023; Ravoux et al., 2023; Karaçaylı et al., 2024; Herrera-Alcantar et al., 2023; Bault et al., 2025; Ramírez-Pérez et al., 2024). Guadalupe was processed using the same set of calibration files and software tags as Fuji, the specprod used for the EDR (DESI Collaboration et al., 2024b). We emphasize, however, that all new or planned analyses of DESI data should use Iron instead of Guadalupe, since Iron contains significantly more data than Guadalupe and was processed with better algorithms and calibration files (see §3.1).
@@
Table 1 summarizes the connection between the public data releases (EDR and DR1) and the spectroscopic productions included in these releases. Unless otherwise noted, in this paper we use DR1 to mean “observations, files, and measurements based on the Iron spectroscopic production,” including all summary statistics and performance metrics. Finally, we point out that some DESI papers, including most of the cosmological analysis papers (e.g., DESI Collaboration et al., 2025b, 2024f) use Year 1 (or Y1) to mean the DR1 or Iron spectroscopic production. So for all intents and purposes, DR1, Y1, and Iron can be used interchangeably to refer to the same underlying set of DESI data products.
@@
We conclude the introduction by outlining the remaining structure of the paper. In §2, we provide an overview of DESI as an instrument (§2.1), describe some of the key DESI target-selection concepts (§2.2), and summarize the observations contained in DR1 and the status of the DESI main survey (§2.3). Next, §3 describes how DESI derives redshifts and spectral classifications from the raw data (§3.1); tabulates and visualizes the number and distribution of unique, confident redshifts in DR1 (§3.2); describes the major DR1 data products and how the data are organized (§3.3); summarizes the large-scale structure catalogs used in the DESI Y1 cosmological analyses (§3.4); and presents the current set of DR1 value-added catalogs (§3.5). Finally, in §4 we describe how the public data can be accessed using two complementary interfaces, and in §5 we summarize the paper.
2 Data Acquisition
@@
In this section we document how DESI targets and observes objects over its deg2 footprint. In §2.1 we describe the characteristics of DESI as a highly multiplexed spectroscopic instrument; in §2.2 we briefly summarize how primary, secondary, and tertiary targets are selected for observations; and in §2.3 we outline DESI survey operations, the process by which DESI carries out its observing program, and summarize the observations contained within DR1.
2.1 Dark Energy Spectroscopic Instrument
@@
2.2 Target Selection
DESI Collaboration et al. (2024a) and Myers et al. (2023a) provide a summary of the selection criteria DESI uses to identify targets for spectroscopic observations, including estimates of the completeness, contamination rate, and technical implementation (building on a significant body of work documented in DESI Collaboration et al. 2016a, and references therein). In addition, individual supporting papers delve into the target-selection algorithms for each specific class of objects targeted by DESI, which we summarize in Table 3 (updated from a prior version of this table presented in Myers et al., 2023a). Here, we briefly describe some key target-selection concepts and terminology used elsewhere in the paper, but defer to the individual papers cited in Table 3 for additional details regarding DESI target selection.
Targets selected as part of the initial DESI design specifications (BGS, ELG, LRG, QSO, and MWS targets; DESI Collaboration et al., 2016a) are typically referred to as primary targets. Primary targets are always photometrically selected from the Legacy Surveys or from Gaia (Gaia Collaboration et al. 2016; see Myers et al. 2023a for details). As mentioned in §1.1, MWS and BGS targets are observed in bright observing conditions while LRG, ELG, and QSO targets are observed in dark conditions (see §2.3.1 and Schlafly et al. 2023) for a quantitative definition of bright and dark observing conditions). In addition, DESI observes stellar backup targets during twilight, or when conditions are too poor for main-survey observations; these backup targets are considered part of the MWS primary program (Dey et al. 2025, in preparation).
@@ 2.2 Target Selection
| Category | Bit NameaaStored as bit-values in the DESI_TARGET, BGS_TARGET and MWS_TARGET columns. Bit-names can be converted to bit-values using the desi_mask, bgs_mask and mws_mask bitmasks (see §2.4 in Myers et al., 2023a). | Description | Reference |
@@
| BGS_WISE | AGN-like BGS target | Juneau et al. (2024a) | |
@@
| Part of the | BACKUP_GIANT_LOPffBACKUP bits are stored in the MWS_TARGET column and mws_mask bitmask. | Candidate Giant Star | Dey et al. 2025 (in prep.) |
@@
In addition to primary targets, DESI also targets secondary and tertiary targets, in order to facilitate bespoke science programs with goals beyond the primary DESI experiment (e.g., Palmese et al., 2021; Darragh-Ford et al., 2023; Dey et al., 2023; Saulder et al., 2023; Yang et al., 2023; Fawcett et al., 2023; Soumagnac et al., 2024; Manser et al., 2024; Byström et al., 2024; Valluri et al., 2025). Secondary targets can be selected using any source of imaging, although the vast majority of these targets have counterparts in the Legacy Surveys. Tertiary targets are similar to secondary targets, but they are observed on dedicated special survey tiles instead of being interleaved with regular targets on normal tiles.
Targets identified for spectroscopic observations are carefully tracked for downstream redshift and large-scale structure catalogs (Ross et al. 2025) using a unique TARGETID (see §3 of Myers et al. 2023a). In addition, we record the provenance of each target using names and values stored in dedicated bit-masks (see §2 of Myers et al. 2023a and Table 3). In Appendix B, C and D we summarize the bits used to track primary, secondary and tertiary targets observed as part of the DESI main survey, respectively; the bits used to track SV target classes are documented in the appendices of the EDR paper, DESI Collaboration et al. (2024b).
2.3 Survey Operations

### v2

@@ Abstract
I Introduction
I.1 The First Year of DESI Data
Elucidating the nature of dark energy and the physical mechanisms responsible for the accelerating expansion of the universe is one of the most important unsolved problems in physics and, arguably, all of science (albrecht06a; weinberg13a). To tackle this question, in 2021 May the Dark Energy Spectroscopic Instrument (DESI) collaboration began a 5-year survey to produce the most detailed three-dimensional map of the universe ever assembled (Levi et al., 2013; DESI Collaboration et al., 2016a, b). By measuring the baryon acoustic oscillation (BAO) feature and redshift space distortions at multiple cosmological epochs, DESI aims to place unprecedented constraints on the equation of state of dark energy, the gravitationally driven growth of large-scale structure, and the sum of the neutrino masses, as well as to explore the observational signatures of primordial inflation (DESI Collaboration et al., 2024a). Concurrently, DESI is carrying out an ambitious survey of stars in the halo of the Milky Way Galaxy in order to constrain the geometry, properties, and accretion history of its stellar halo, disk, and dark-matter halo (Cooper et al., 2023). Notably, DESI has the distinction of being the first Stage-IV dark-energy experiment to begin science operations (albrecht06a).
@@ I.1 The First Year of DESI Data
DESI identifies its primary spectroscopic targets using 14,000 deg2 of broadband optical and mid-infrared photometry from Data Release 9 (DR9) of the DESI Legacy Imaging Surveys111https://www.legacysurvey.org/dr9 (Zou et al. 2017; Dey et al. 2019; Moustakas et al. 2023; Myers et al. 2023; hereafter, the Legacy Surveys), and, in some cases, using Gaia stellar photometry (gaia-collaboration16a). Specifically, DESI targets five broad classes of objects scaffolded in redshift: Milky Way survey (MWS) and backup program stars (Cooper et al., 2023; dey25a); bright galaxy survey (BGS) galaxies (; Ruiz-Macias et al., 2020; Hahn et al., 2023); luminous red galaxies (LRGs, ; Zhou et al., 2020, 2023); emission-line galaxies (ELGs, ; Raichoor et al., 2020, 2023); and QSOs (; Yèche et al., 2020; Chaussidon et al., 2023). For its cosmological analyses, DESI further differentiates QSO targets into “tracer” QSOs at lower redshift, , from Ly forest QSOs at , because above DESI uses the Ly forest as an independent probe of the matter-density field (DESI Collaboration et al., 2016a).
@@
After first light and a brief commissioning period, in 2020 December DESI began a five-month survey validation (SV) program which was designed to test the performance of the instrument and all its subsystems, and to validate its target-selection algorithms, data-reduction pipeline, and scientific deliverables (DESI Collaboration et al., 2024a). SV consisted of three successive phases: Target Selection Validation (SV1); Operations Development (SV2); and the One-Percent Survey (SV3), the last of which covered roughly 1% of the final 14,000 deg2 DESI footprint but to higher target completeness (DESI Collaboration et al., 2024a). These observations, particularly the One-Percent Survey data, resulted in a flurry of scientific activity222https://data.desi.lbl.gov/doc/papers/edr as well as DESI’s first measurement of the BAO peak in the BGS, LRG, ELG, and QSO 2-point correlation functions (Moon et al., 2023) and in the auto-correlation function of the Ly forest (gordon23a). Subsequently, on 2021 May 14, DESI launched its 5-year main survey.333In detail, a small number of SV observations continued until 2021 June 10, but those data were acquired independently of the main survey and are included in the SV dataset (DESI Collaboration et al., 2024b).
@@
All the data obtained during SV were first made public on 2023 June 13 as part of the DESI Early Data Release (EDR; DESI Collaboration et al., 2024a, b).444https://data.desi.lbl.gov/doc/releases/edr Here, we present DESI Data Release 1 (DR1), which includes all the data obtained by DESI during its first 13 months of science operations (2021 May 14 through 2022 June 13) as well as a uniform reprocessing of all the SV data. Overall, DESI has been efficient, running well-ahead of schedule and occasionally obtaining redshifts for more than one million unique targets in a single month (Schlafly et al., 2023). In the first year of the main survey, DESI has measured confident redshifts for approximately 18.7M unique targets over more than deg2, including 13.1M galaxies, 1.6M quasars, and 4M stars, making DESI the largest extragalactic spectroscopic redshift survey ever conducted. For comparison, all five generations of the Sloan Digital Sky Survey (SDSS I/II, III, IV, and V; york00a; dawson13a; majewski17a; blanton17a; kollmeier17a), spanning 18 public data releases and approximately 25 years of operations,555https://www.sdss.org have cumulatively resulted in approximately 4M unique extragalactic spectra, making DESI DR1 nearly a factor of four larger than all previous SDSS programs combined.
@@
In §I.2 we outline the organization and contents of this paper and describe several important terms and concepts which will help orient readers interested in using DESI DR1. Complementarily, we recommend that readers interested in the cosmological results from DESI DR1 begin at the DESI portal666https://data.desi.lbl.gov/doc/papers/#year-1-cosmology-results for a high-level overview of the results before consulting the individual key papers (DESI Collaboration et al., 2024d, f, 2025b, e, 2025a, c).
I.2 A High-Level Guide to DESI and Data Release 1
@@
All DESI observations are made in the context of a survey and a program (Schlafly et al., 2023). As introduced in §I.1, DESI’s primary, 5-year scientific program is its main survey, which was preceded by three independent, successive surveys (SV1, SV2, and SV3), collectively called SV. In addition, an object can be observed as part of the special survey, which includes a miscellaneous assortment of targets and observations which are kept separate from the main survey (see §II.3.3).
Within each survey, objects are observed in one of three possible programs according to the acceptable lunar phase and on-sky observing conditions (see §II.3.1): bright, dark, or backup.777A fourth program name, other, is occassionally used for ad hoc or bespoke observations which do not easily fit into the three standard program names. By design, LRG, ELG, and QSO targets are assigned to the dark program and BGS and MWS targets are assigned to the bright program; meanwhile, the backup program consists entirely of stellar (Milky Way) targets, and special (secondary and tertiary targets) can be assigned to any program (see §II.2; Schlafly et al., 2023; Myers et al., 2023; DESI Collaboration et al., 2024a). In general, objects within each survey/program combination are always treated independently, from target selection (§II.2) all the way through to spectral coaddition and redshift estimation (§III.1); in other words, the same astrophysical object can appear in two (or more) survey/program combinations. This strict separation of targets within each survey and program is a critical part of the large-scale structure pipeline which enables DESI to make precise cosmological measurements (Ross et al., 2025).
@@
| ProductionaaWithin the DESI data model, a spectroscopic production is referred to as a specprod. Each specprod corresponds to a uniform processing of the data using a well-defined and reproducible set of input files and software tags. | Stack | SurveysbbSee §II.3 for a description of the SV and main surveys. | RedshiftsccApproximate total number of unique redshifts summed over all surveys and programs. | Date |
@@
The primary spectroscopic production for DR1 is Iron, named after the Iron Mountain in Utah, USA; the Iron spectroscopic production includes data from the first 13 months of the DESI main survey, as well as a uniform reprocessing of the SV data which was previously released in the EDR (DESI Collaboration et al., 2024b). DR1 also includes a supplemental or ancillary spectroscopic production called Guadalupe (after the Guadalupe Mountains in Texas, USA), which includes just the first two months of the DESI main survey data (see Appendix A). We include Guadalupe in DR1 because it was used in some early DESI analyses (e.g., Moon et al., 2023; gordon23a; ravoux23a; karacayli24a; herrera-alcantar23a; Bault et al., 2025; ramirez-perez24a). Guadalupe was processed using the same set of calibration files and software tags as Fuji, the specprod used for the EDR (DESI Collaboration et al., 2024b). We emphasize, however, that all new or planned analyses of DESI data should use Iron instead of Guadalupe, since Iron contains significantly more data than Guadalupe and was processed with better algorithms and calibration files (see §III.1).
@@
Table 1 summarizes the connection between the public data releases (EDR and DR1) and the spectroscopic productions included in these releases. Unless otherwise noted, in this paper we use DR1 to mean “observations, files, and measurements based on the Iron spectroscopic production,” including all summary statistics and performance metrics. Finally, we point out that some DESI papers, including most of the cosmological analysis papers (e.g., DESI Collaboration et al., 2025a, 2024c) use Year 1 (or Y1) to mean the DR1 or Iron spectroscopic production. So for all intents and purposes, DR1, Y1, and Iron can be used interchangeably to refer to the same underlying set of DESI data products.
@@
We conclude the introduction by outlining the remaining structure of the paper. In §II, we provide an overview of DESI as an instrument (§II.1), describe some of the key DESI target-selection concepts (§II.2), and summarize the observations contained in DR1 and the status of the DESI main survey (§II.3). Next, §III describes how DESI derives redshifts and spectral classifications from the raw data (§III.1); tabulates and visualizes the number and distribution of unique, confident redshifts in DR1 (§III.2); describes the major DR1 data products and how the data are organized (§III.3); summarizes the large-scale structure catalogs used in the DESI Y1 cosmological analyses (§III.4); and presents the current set of DR1 value-added catalogs (§III.5). Finally, in §IV we describe how the public data can be accessed using two complementary interfaces, and in §V we summarize the paper.
II Data Acquisition
@@
In this section we document how DESI targets and observes objects over its deg2 footprint. In §II.1 we describe the characteristics of DESI as a highly multiplexed spectroscopic instrument; in §II.2 we briefly summarize how primary, secondary, and tertiary targets are selected for observations; and in §II.3 we outline DESI survey operations, the process by which DESI carries out its observing program, and summarize the observations contained within DR1.
II.1 Dark Energy Spectroscopic Instrument
@@
II.2 Target Selection
DESI Collaboration et al. (2024a) and Myers et al. (2023) provide a summary of the selection criteria DESI uses to identify targets for spectroscopic observations, including estimates of the completeness, contamination rate, and technical implementation (building on a significant body of work documented in DESI Collaboration et al. 2016a, and references therein). In addition, individual supporting papers delve into the target-selection algorithms for each specific class of objects targeted by DESI, which we summarize in Table 3 (updated from a prior version of this table presented in Myers et al., 2023). Here, we briefly describe some key target-selection concepts and terminology used elsewhere in the paper, but defer to the individual papers cited in Table 3 for additional details regarding DESI target selection.
Targets selected as part of the initial DESI design specifications (BGS, ELG, LRG, QSO, and MWS targets; DESI Collaboration et al., 2016a) are typically referred to as primary targets. Primary targets are always photometrically selected from the Legacy Surveys or from Gaia (gaia-collaboration16a; see Myers et al. 2023 for details). The photometric selection criteria include carefully tuned color-cuts designed to maximize the number density and redshift efficiency of each target class, in addition to the following magnitude limits: mag (BGS_BRIGHT); mag (BGS_FAINT); mag (LRGs); mag (ELGs); and mag (QSOs). In Table 3 we list the references which justify and document the complete set of magnitude and color-selection criteria adopted for each class (see also Myers et al. 2023).
MWS and BGS targets are observed in bright observing conditions while LRG, ELG, and QSO targets are observed in dark conditions (see §II.3.1 and Schlafly et al. 2023 for a quantitative definition of bright and dark observing conditions). In addition, DESI observes stellar backup targets during twilight, or when conditions are too poor for main-survey observations; these backup targets are considered part of the MWS primary program (Dey et al. 2025, in preparation).
@@ II.2 Target Selection
| Category | Bit NameaaStored as bit-values in the DESI_TARGET, BGS_TARGET and MWS_TARGET columns. Bit-names can be converted to bit-values using the desi_mask, bgs_mask and mws_mask bitmasks (see §2.4 in Myers et al., 2023). | Description | Reference |
@@
| BGS_WISE | AGN-like BGS target | juneau25a | |
@@
| Part of the | BACKUP_GIANT_LOPffBACKUP bits are stored in the MWS_TARGET column and mws_mask bitmask. | Candidate Giant Star | dey25a |
@@
In addition to primary targets, DESI also targets secondary and tertiary targets, in order to facilitate bespoke science programs with goals beyond the primary DESI experiment (e.g., palmese21a; darragh-ford23a; dey23a; saulder23a; yang23a; fawcett23a; soumagnac24a; manser24a; bystrom25a; valluri25a). Secondary targets can be selected using any source of imaging, although the vast majority of these targets have counterparts in the Legacy Surveys. Tertiary targets are similar to secondary targets, but they are observed on dedicated special survey tiles instead of being interleaved with regular targets on normal tiles.
Targets identified for spectroscopic observations are carefully tracked for downstream redshift and large-scale structure catalogs (Ross et al. 2025) using a unique TARGETID derived from its sky position and the imaging data release from which it was selected; consequently, cross-matching across observations or catalogs is performed exclusively by TARGETID rather than by sky position, ensuring unambiguous identification of the same astrophysical object across all DESI observations (see §3 of Myers et al. 2023 for details on the TARGETID construction).
In addition, we record the provenance of each target using names and values stored in dedicated bit-masks (see §2 of Myers et al. 2023 and Table 3). In Appendix B, C and D we summarize the bits used to track primary, secondary and tertiary targets observed as part of the DESI main survey, respectively; the bits used to track SV target classes are documented in the appendices of the EDR paper, DESI Collaboration et al. (2024b).
II.3 Survey Operations